OpenStudy (anonymous):

Which of the following does not represent the set of integers from 1 to 9?

OpenStudy (anonymous):

Which of the following does not represent the set of integers from 1 to 9?

jimthompson5910 (jim_thompson5910):

You're missing "the following" choices

jimthompson5910 (jim_thompson5910):

Are you given answer choices?

OpenStudy (anonymous):

{integers from 1 to 9} {1, 2, 3, 4, 5, 6, 7, 8, 9} {1, 2, 3, …, 7, 8, 9} {integers greater than 1}

jimthompson5910 (jim_thompson5910):

{integers from 1 to 9} = {1,2,3,4,5,6,7,8,9}

jimthompson5910 (jim_thompson5910):

so choices A and B are the same

jimthompson5910 (jim_thompson5910):

so is C

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

Which of the following represents a set of integers greater than zero? {1, 2, 3, 4, 5} {1, 2, 3, …, 98, 99, 100} {all positive numbers} {1, 2, 3, 4, 5, …}

jimthompson5910 (jim_thompson5910):

which one do you think it is

OpenStudy (anonymous):

c

OpenStudy (anonymous):

or b

jimthompson5910 (jim_thompson5910):

good, D works as well

OpenStudy (anonymous):

oohhh lol thnanks

jimthompson5910 (jim_thompson5910):

B does not since 101 is an integer greater than 0, but it's not in set B

OpenStudy (anonymous):

oohh i get that onne now thanks

jimthompson5910 (jim_thompson5910):

does it say anything more about this set?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

hmm odd

OpenStudy (anonymous):

i dont understand this everytime i try i get it wrong :/ Part 1: Write your own real world situation that can be organized into a Venn diagram with two overlapping circles. Part 2: Provide the appropriate information, specific to your created problem from Part 1, for the generic Venn diagram below. What is the title of your Venn diagram? What is the title of Circle A? What is the title of Circle B? What are the values of v, x, y, and z?

jimthompson5910 (jim_thompson5910):

think of two sets

jimthompson5910 (jim_thompson5910):

ex: set of people who like ice cream and set of people who like cookies

OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = x + 4 and passes through the point (−6, 3). y = x − 1 y = x + 3 y = −x − 1 y = −x + 3

OpenStudy (anonymous):

and thanks for the example

jimthompson5910 (jim_thompson5910):

What is the slope of y = x + 4

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

no, think of y = x+4 as y = 1x+4

OpenStudy (anonymous):

is it 1?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

or 1/1

jimthompson5910 (jim_thompson5910):

To find the perpendicular slope, flip the fraction and change the sign to get -1/1

jimthompson5910 (jim_thompson5910):

which simplifies to -1

OpenStudy (anonymous):

so its a? like do you go by y intercept?

jimthompson5910 (jim_thompson5910):

now use y = mx+b with m = -1, x = -6, and y = 3 y = mx+b 3 = -1*(-6)+b Now solve for b

OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x − y = −3 and passes through the point (4, −2). y = −x − 6 y = −x − 14 y = 3x − 6 y = 3x − 14

OpenStudy (anonymous):

its cutting off most of the question idk idk idk im getting stressed

jimthompson5910 (jim_thompson5910):

3x − y = C 3(4) − (-2) = C keep going to find C tell me what you get

jimthompson5910 (jim_thompson5910):

Once you find C, you need to rewrite the equation in slope intercept form

jimthompson5910 (jim_thompson5910):

Does that make sense?